# Position: Meaning Is Not A Metric: Using LLMs to make cultural context legible at scale

### v1

@@ 4.4 What $\neq$ How Much
At first glance, creating a mapping between these two facets of meaning may seem like a straightforward empirical problem; however, this is not the case [15, 35]. For instance, it might seem that the solution would simply be to gather enough human-generated data in which people rate experiences as more or less meaningful. With enough features coded into the data, it could be determined which experiences people find meaningful—and those could be supported through technological interventions. However, this kind of approach would have difficulty accounting for the ambiguity in asking participants about meaning versus happiness [69], the radical flexibility of the interpretive process underlying judgments of meaning [52], the difficulties of codifying context as a feature set [113, 44, 85], that the same events can be perceived as more or less meaning depending on their narrative framing [83], the discrepancies between self-report and revealed preference [63], among many other difficulties in combining cultural and psychological perspectives on meaning [11, 98]. It therefore remains an open question how best to develop mappings from qualitative interpretations about the cultural content of meaning to quantitative estimates of its psychological magnitude.
@@ 4.5 Meaning Is Made, Not Found
Recent psychological evidence shows that people’s judgments about meaning can change based on the narrative framing of their experiences [83]: the same events elicit different ratings of meaningfulness when they occur in the context of different stories. This suggests that not only will LLMs be useful for measuring meaning—but they will become participants in the meaning-making process. For example, they may influence people’s perceptions about the meaningfulness of their own experiences through guided journaling [68, 46], personalized personas [18], or psychotherapy [67, 45]. Furthermore, how LLMs categorize people as having a shared contextual frame may have crucial downstream effects on their interactions with others [16]. Overall, this is consistent with what humanists and social scientists call the “reflexivity” of meaning: that the way we measure, codify, or understand meaning has a direct impact on the meaning-making process itself [31, 110, 8].
@@ 5 Discussion
In this paper, we have argued that LLMs can be used to make human meaning legible at an unprecedented scale. We contend that the primary obstacle to this effort is that human meaning can only be encoded via thick description—but, historically, the only way to reach legibility at scale is via thin description. We propose that LLMs can be used to make thick descriptions a useful representational strategy in large-scale sociotechnical systems. A crux of this is about adequately representing cultural context. Accordingly, there are useful gradations of thickness between descriptions that are fully thick and fully thin—and therefore many ways to usefully make progress on this problem. These gradations exist through relaxation of the mandate for standardization, mainly by finding ways to accommodate the inherent heterogeneity of socially-situated behavior.
@@ 5.1 Better Metrics Are Not Enough
Our proposal shares a premise with other calls to use AI in reshaping our society’s sociotechnical infrastructure (such as recommender systems) to reflect a broader sense of human well-being [97, 95, 42]. However, these efforts are primarily concerned with identifying which proxy metrics will most effectively capture human well-being. Our position is that all metrics will inevitably succumb to proxy failure [43] and what is needed instead is an entirely new representational format (i.e., thick description). To illustrate, this is why scholars who are concerned with meaning in the humanities and social sciences write papers full of long verbal exposition rather than ledgers of numbers [27, 28, 29]. Our proposal is similarly aligned to the nascent field of representation engineering [115], which places the representation of information (rather than its processing) at the core of technical development.
@@ 5.3 This Is A Crucial Application Of Approaches To Scaling Qualitative Methods
Social scientists have begun exploring a range of efforts to scale qualitative methods (including thick description) using LLMs and other forms of generative AI [79, 23, 88, 49, 65, 107, 81]. Some offer thick description (or other social science approaches) as a method for evaluating AI systems [79, 107, 88]. A notable case is Qadri and colleagues (2025), who argue that “thick evaluations” are necessary to measure whether AI can represent key aspects of minority cultures [79]. While our position is aligned with their work in many ways, we argue that thickness is needed to represent all cultures (not just minorities) and that it should be used as a representational scheme itself and not just for evaluation. Other efforts have used LLMs to conduct qualitative analysis on large corpora of culturally-situated text [23, 81, 49, 65]; building on this work will be foundational toward ensuring the reliability of deploying thick description at scale. Especially promising work in this vein uses qualitative methods in not just in data analysis but in the generation of data as well [62].
@@ 5.4 Alternative Views, Risks, and Points of Failure
This is going to lead to bad outcomes because you will fail at robustly representing meaning; like social media, you will start off trying to measure “what really matters” but ultimately end up with something much more narrow insidiously masquerading as the deeper thing. As implied in our section presenting five challenges, a lot can go wrong when trying to represent meaning. It is easy to overextend a given generalization beyond where it would be useful, accurate, or beneficial. It is difficult to convey the essence of culture in a way that does not trivialize it [113, 44]. One way our proposal could go wrong is by claiming to liberate a sociotechnical system from “enforced standardization”—only to make such standardization harder to discern. The risk is overgeneralization: applying one cultural standard for meaning in a context (or to an individual) where it does not belong.
@@
You claim to draw on humanities and social science frameworks. But thick description is only one such framework; others paint a much more complicated picture. For example, Haraway (1988) argues that all knowledge is intrinsically partial, embodied, and dependent on specific historical and social contexts [33]. Thus, when the scientific community assumes that the perspective of its work is neutral and objective, it is simply imposing its own partial, contextually-bound perspective—usually a WEIRD one [39]—on an authoritative body of knowledge without having to acknowledge the imposition. According to Haraway, this paper’s proposal could potentially provide a more powerful mechanism for imposing a singular way of seeing things on a diverse cultural ecosystem. Likewise, Barad (2007) argues that the act of making something legible within a sociotechnical system does not just observe it in a neutral sense, but in fact transforms it [4]. For example, when a recommender system measures user preferences, it is not just recording pre-existing preferences; it is creating the categories that make certain preferences possible, likely, or even inexpressible. According to Barad, this paper’s proposal could potentially exert a corrupting influence on the original cultural phenomenon itself, pressuring it to fit into the categorical box that renders it legible within the sociotechnical system.
@@
6 Acknowledgements
Thanks to Danny Oppenheimer, who offered in-depth commentary, and to the attendees of the October 2024 Pittsburgh “Doing AI Differently” workshop, whose discussions helped spark the ideas that set this paper in motion.

### v2

@@ 4.4 What $\neq$ How Much
At first glance, creating a mapping between these two facets of meaning may seem like a straightforward empirical problem; however, this is not the case [15, 36]. For instance, it might seem that the solution would simply be to gather enough human-generated data in which people rate experiences as more or less meaningful. With enough features coded into the data, it could be determined which experiences people find meaningful—and those could be supported through technological interventions. However, this kind of approach would have difficulty accounting for the ambiguity in asking participants about meaning versus happiness [77], the radical flexibility of the interpretive process underlying judgments of meaning [55], the difficulties of codifying context as a feature set [122, 46, 92], the observation that the same events can be perceived as more or less meaningful depending on their narrative framing [90], the discrepancies between self-report and revealed preference [70], among many other difficulties in combining cultural and psychological perspectives on meaning [11, 107]. It therefore remains an open question how best to develop mappings from qualitative interpretations about the cultural content of meaning to quantitative estimates of its psychological magnitude.
@@ 4.5 Meaning Is Made, Not Found
Recent psychological evidence shows that people’s judgments about meaning can change based on the narrative framing of their experiences [90]: the same events elicit different ratings of meaningfulness when they occur in the context of different stories. This suggests that not only will LLMs be useful for measuring meaning—but they will become participants in the meaning-making process [57]. For example, they may influence people’s perceptions about the meaningfulness of their own experiences through guided journaling [75, 48], personalized personas [19], psychotherapy [74, 47], or even the kind of content they create and share online [56]. Furthermore, how LLMs categorize people as having a shared contextual frame may have crucial downstream effects on their interactions with others [16]. Overall, this is consistent with what humanists and social scientists call the “reflexivity” of meaning: that the way we measure, codify, or understand meaning has a direct impact on the meaning-making process itself [32, 119, 8].
5 Call to Action
What does this look like in practice? The position offered in the paper articulates a space of potential approaches, not a single one. With that in mind, we can point to what we consider to be an especially promising direction: domain-agnostic qualitative coding. In the social sciences, the standard procedure for translating between thicker qualitative representations and thinner quantitative ones is called “qualitative coding” [94, 106]. Suppose you are social scientist studying a dataset from social media, looking at signs of psychological distress or mental instability. As a researcher, you would develop a “code” which thoroughly defines the qualitative aspects of the signal you are trying to identify, along with a numerical scheme for assessing the degree to which that signal is present (e.g., you might have a paragraph that describes how people use punctuation differently when they are enraged and would then score the presence of that behavior as 0 [“not present”] to 3 [“highly present”] for a given social media post). This gives a concrete example of what it looks like to aggregate thick qualitative descriptions into a numerical format.
The limitations of qualitative coding for our proposal is that each application requires a bespoke code. There is no secret guide given to students on the first day of Sociology 101 which holds the formula for one-code-to-rule-them-all. For each study, a new code must be designed. This typically is done by someone with a high degree of expertise in their field. The execution of that code does not need to be done by someone with a PhD, only someone trained in understanding and employing the code (e.g., an undergraduate research assistant). Our proposal is that AI could partially automate not only the application of coding schemes—which is already of great interest in the social sciences [68, 106, 43, 20]—but also the generation of bespoke codes in real time. This is what we mean by this approach to scaling thick description being “domain-agnostic”: the LLM would be designed to create an appropriate code for a given situation and identify pertinent themes at hand before applying it. Crucially, a mechanism would be needed to cross reference these with human judgments; for example, by presenting professional human coders with a subset of the same stimuli—in an effort to create a collaborative dynamic for producing codes at scale, rather than pure automation.
This proposal is not something that can be accomplished in a single research project. There are many ways to implement different versions of this approach. There are many ways that it could go wrong or be unreliable. There are many ways to aggregate qualitative representations into numerical ones. There are many ways to incorporate human feedback or oversight into this process. This multiple-realizability is, ultimately, why this proposal is presented as a position and not a single research paper. We propose that a concrete route to “using LLMs to make culture context legible at scale” is to begin addressing these questions in an overall effort to develop technical approaches for domain-agnostic qualitative coding at scale.
@@ 6 Discussion
In this paper, we have argued that LLMs can be used to make human meaning legible at an unprecedented scale. We contend that the primary obstacle to this effort is that human meaning can only be encoded via thick description—but, historically, the only way to reach legibility at scale is via thin description. We propose that LLMs can be used to make thick descriptions a useful representational strategy in large-scale sociotechnical systems. There are useful gradations of thickness between descriptions that are fully thick and fully thin—and therefore many ways to usefully make progress on this problem. These gradations exist through relaxation of the mandate for standardization, mainly by finding ways to accommodate the inherent heterogeneity of socially-situated behavior.
@@ 6.1 Better Metrics Are Not Enough
Our proposal shares a premise with other calls to use AI in reshaping our society’s sociotechnical infrastructure (such as recommender systems) to reflect a broader sense of human well-being [103, 104, 44, 62]. However, these efforts are primarily concerned with identifying which proxy metrics will most effectively capture human well-being. Our position is that all metrics will inevitably succumb to proxy failure [45] and what is needed instead is an entirely new representational format (i.e., thick description). To illustrate, this is why scholars who are concerned with meaning in the humanities and social sciences write papers full of long verbal exposition rather than ledgers of numbers [28, 29, 30]. Our proposal is similarly aligned to the nascent field of representation engineering [124], which places the representation of information (rather than its processing) at the core of technical development.
@@ 6.3 This Is A Crucial Application Of Approaches To Scaling Qualitative Methods
Social scientists have begun exploring a range of efforts to scale qualitative methods (including thick description) using LLMs and other forms of generative AI [86, 24, 95, 51, 72, 116, 88]. Some offer thick description (or other social science approaches) as a method for evaluating AI systems [86, 116, 95, 62]. A notable case is Qadri and colleagues, who argue that “thick evaluations” are necessary to measure whether AI can represent key aspects of minority cultures [86]. Likewise, Lowe and colleagues articulate a vision for “full-stack alignment” based on “thick” models of value [62]. Other efforts have used LLMs to conduct qualitative analysis on large corpora of culturally-situated text [24, 88, 51, 72]; building on this work will be foundational toward ensuring the reliability of deploying thick description at scale. Especially promising work in this vein uses qualitative methods in not just in data analysis but in the generation of data as well [69].
@@ 6.4 Alternative Views, Risks, and Points of Failure
This is going to lead to bad outcomes because you will fail at robustly representing meaning; like social media, you will start off trying to measure “what really matters” but ultimately end up with something much more narrow insidiously masquerading as the deeper thing. As implied in our section presenting five challenges, a lot can go wrong when trying to represent meaning. It is easy to overextend a given generalization beyond where it would be useful, accurate, or beneficial. It is difficult to convey the essence of culture in a way that does not trivialize it [122, 46]. One way our proposal could go wrong is by claiming to liberate a sociotechnical system from “enforced standardization”—only to make such standardization harder to discern. The risk is overgeneralization: applying one cultural standard for meaning in a context where it does not belong.
@@
You claim to draw on humanities and social science frameworks. But thick description is only one such framework; others paint a much more complicated picture. For example, Haraway argues that all knowledge is intrinsically partial, embodied, and dependent on specific historical and social contexts [34]. Thus, when the scientific community assumes that the perspective of its work is neutral and objective, it is simply imposing its own partial, contextually-bound perspective—usually a WEIRD one [40]—on an authoritative body of knowledge without having to acknowledge the imposition. According to Haraway, this paper’s proposal could potentially provide a more powerful mechanism for imposing a singular way of seeing things on a diverse cultural ecosystem. Likewise, Barad argues that the act of making something legible within a sociotechnical system does not just observe it in a neutral sense, but in fact transforms it [4]. For example, when a recommender system measures user preferences, it is not just recording pre-existing preferences; it is creating the categories that make certain preferences possible, likely, or even inexpressible. According to Barad, this paper’s proposal could potentially exert a corrupting influence on the original cultural phenomenon itself, pressuring it to fit into the categorical box that renders it legible within the sociotechnical system.